Assume you buy a bond with the following features Bond maturity = 4 Coupon Rate = 5.00% Face Value = $1,000 Annual Coupons When you buy the bond the market interest rate = 6.00% Immediately after you buy the bond the interest rate changes to 5.50% What is the "reinvestment" effect in year 3 ? Group of answer choices a) -$0.78 b) -$0.80 c) $0.80 d) $0.78
Answer #1

Answer - Option (a) i.e. -0.78

Explanation

Reinvestment risk is the risk referred to as risk that bond return is invested at a lower rate in the future.

FV of Coupon at the end of 3rd year, reinvested at market rate i.e. 5.5% p.a.

= (1000*5%)*[(1+5.5%)^3-1]/5.5%

= (50)*[1.174241375-1]/5.5%

= 158.40

FV of Coupon at the end of 3rd year, if reinvested @ 6% p.a.

= (1000*5%)*[(1+6%)^3-1]/6%

= (50)*[1.1191016-1]/6%

= 159.18

Hence the reinvestment effect = 158.40 - 159.18 = -0.78